28 February 2008

RHIC Running:  Maintenance Day yesterday went well.  PHENIX got their MuTR trigger work done.  Most of the PP2PP development work got done.  Booster and AGS work got done, including the target replacement for the AGS Polarimeter and the polarized proton source maintenance.  Recovery from Maintenance also went well, with a 12 bunch store and a physics store by 8:00PM.

The APEX work on beta*=60cm development was stymied by loading two incorrect ramp files.  pp83 instead of pp83lowbeta.  They need a couple of good ramps with the correct ramp file in order for collimation work and tuning to be completed. 

The STAR ZDC rates are higher than the PHENIX ZDC rates now.  This is a change for this run, and it is not understood .  Wolfram is concerned that we are using the wrong cross sections for calculating emittances.  Previous running used 41 and 50 nb for PHENIX and STAR, whereas this year it appears to be 31/29 nb.  We need to use all our tools to study the emittance and come to a consensus. 

STAR requests an half hour access to work on their data acquisition system to improve stability.  This will occur immediately after the beam is dumped at 13:00.  Prior to dumping the beam, there will be a polarization measurement and a 10 minute study for the stochastic cooling.


Scheduling Physicist: Mike Sivertz Last updated:
Email: sivertz@bnl.gov
Office Phone: 631-344-6102
Digital Pager: 7528 or 4300
Home Phone: 631-878-9020